OpenStudy (anonymous):

If sin(α) = 24/25 where 0 < α < π/2 and cos(β) = 21/29 where 3π/2 < β < 2π. sin(α) = 24/25 sin(β) = -7/25 cos(α)= 20/29 cos(β)= 21/29 how do you do tan(α − β)? i already know the formula but i can't get the right answer

OpenStudy (kc_kennylau):

well you used the wrong formula

OpenStudy (kc_kennylau):

the formula is \(\sin^2\theta+\cos^2\theta=1\)

OpenStudy (anonymous):

i think i used the right one

OpenStudy (kc_kennylau):

\[\sin\alpha=\dfrac{24}{25}\]\[\cos\alpha=\dfrac7{25}\] \[\cos\beta=\dfrac{21}{29}\]\[\sin\beta=\dfrac{20}{29}\]

OpenStudy (kc_kennylau):

how can \(\cos\beta\) be positive if \(\dfrac{3\pi}2<\beta<2\pi\)?

OpenStudy (anonymous):

oh. sorry can you help me solve it.

OpenStudy (kc_kennylau):

Oh sorry let me do that again

OpenStudy (anonymous):

okay

OpenStudy (kc_kennylau):

\[\sin\alpha=\frac{24}{25}\]\[cos\alpha=\sqrt{1-\sin^2\alpha}=\frac7{25}\] \[\cos\beta=\frac{21}{29}\]\[\sin\beta=-\sqrt{1-\cos^2\alpha}=-\frac{20}{29}\]

OpenStudy (kc_kennylau):

formula: \(\tan(\alpha-\beta)=\dfrac{\tan\alpha-\tan\beta}{1+\tan\alpha\tan\beta}\)

OpenStudy (kc_kennylau):

\[\tan\alpha=\frac{\sin\alpha}{\cos\alpha}=\frac{24}7\] Can you work out \(\tan\beta\) and plug that into the formula? :)

OpenStudy (anonymous):

no i'm really having a hard time right now figuring it out

OpenStudy (kc_kennylau):

\(\tan\beta\) is \(\dfrac{\sin\beta}{\cos\beta}\) now can you work out the value of it?

OpenStudy (anonymous):

-20/21 right?

OpenStudy (kc_kennylau):

yep :)